The average height of students in a freshman class of a certain school has been 158.55 cm with a population standard deviation of 8.8 cm. Is there a reason to believe that there has been a change in the average height if a random sample of 51 students in the present freshman class has an average height of 154.25 cm? Use a 0.1 level of significance.






What are the given? Write only the number. :




population mean:Blank 1 cm




population standard deviation:Blank 2 cm




sample size:Blank 3




sample mean:Blank 4 cm




level of significance:Blank 5






What are the critical values? Write the positive critical value first then the negative.




z:Blank 6 andBlank 7






What is the value of the calculated z? Round your answer to the nearest hundredths.




z:Blank 8

1
Expert's answer
2022-05-18T16:28:55-0400

population mean: "158.55" cm

population standard deviation: "8.8" cm

sample size:"51"

sample mean: "154.25" cm

level of significance: "0.1"

What is the critical value?

"z_{c1}=-1.6449, z_{c2}=1.6449"

What is the value of the calculated z?


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{154.25-158.55}{8.8/\sqrt{51}}\approx-3.49"

The following null and alternative hypotheses need to be tested:

"H_0:\mu=158.55"

"H_1:\mu\not=158.55"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Since it is observed that "z=-3.49<-1.6449= z_{c1}," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(z<-3.49)=0.000483," and since "p=0.000483<0.1=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 158.55, at the "\alpha = 0.1" significance level.